OpenStudy (anonymous):

Simplify each expression. Show all work, step-by-step. (123) ^ -2/3

OpenStudy (australopithecus):

\[123^{-\frac{2}{3}}=\frac{1}{123^{\frac{2}{3}}}\]

OpenStudy (australopithecus):

\[= \frac{1}{\sqrt[3]{123^{2}}}\]

OpenStudy (anonymous):

Thats the answer? @Australopithecus

OpenStudy (australopithecus):

no, but it should help you to simplify it

OpenStudy (anonymous):

how do i go from there

OpenStudy (anonymous):

thats supposed to say 125 sorry not 123

OpenStudy (anonymous):

@diysohail help

OpenStudy (anonymous):

hi,attached.

OpenStudy (anonymous):

the last answer(1/25) can be said like 25^-1